OpenStudy (anonymous):

Coral created a painting that was 4 1/3 ft long and 1 1/5 ft wide. What was the area of Coral's painting?

Mehek (mehek14):

\(\dfrac{13}{3}*\dfrac{6}{5} = \dfrac{78}{15} = 5 \dfrac{1}{2}\)

OpenStudy (anonymous):

thats not part of the answer choice

Mehek (mehek14):

what do you have for answer choices?

Mehek (mehek14):

\(5\dfrac{1}{5}\)

OpenStudy (anonymous):

ok thats better

Mehek (mehek14):

yes just saw my mistake

OpenStudy (anonymous):

can i ask another question

Mehek (mehek14):

yes

OpenStudy (anonymous):

Chico walks 3/5 mile to the library. Nikki walks 1/4 the distance that Chico walks to the library. How far does Nikki walk?